Current P-Way Works 2017.
The P-Way department of the NYMR is based at Newbridge Yard in Pickering. Regular Sunday working takes place most weekends. Attendance is generally from 09.00 to 17.00 unless previously announced.
Work is varied ranging from Track Patrol work to Bridge Rebuilding with a bit of Track Maintenance or Renewal as well. What better way to view the trains in the scenery than from trackside while working for the NYMR!
There are also regular Plant and Rolling stock repairs/maintenance activities and track material preparation at Newbridge on some Sundays, Tuesdays and Fridays.
Volunteers must hold a valid PTS certificate, this can be arranged through P Way Supervisor M Cannings.
If you would like to be involved, please use the contact page to gain information.
Details for 2017 below
29th Dec 2016 –
Bryan and Nick in attendance today.
Remaining pre-drilled holes at north end now fitted with bolts.
External welding seams and remaining rivet residue ground flush. The large steel end plates were then cleaned off and given an external coat of red metal primer.
During this work, we noticed that the naughty Christmas pixies must have been in the depot over the holiday , as we found a socket wedged firmly over a newly installed nut and behind a rivet, in a very inaccessible place.
It took two screwdrivers, a hammer and two of us – sometimes hung upside down ( or nearly) to knock it out.
Snow plough 18
Dressing new interior tongue and groove boarding, easing the sticking east side door and burning off and sanding blistered paint on the plough end.
Mon 3rd Jan –
Dismantling of the F19`s at MP14 will start today once ECS movements have taken place, with all materials being lifted out by the end of the week. Relaying will start immediatly after, and then will be followed by the rerailing of the section in AD sleepers. If all goes well I hope the track to be complete and ready for ballasting by Sunday 22nd of January. The job should proceed quickly due to the fact that the G44`s have already been prefitted with clips , pads and insulators. At this stage I an not planning to work on Saturdays. After this the rails will be loaded up and disposed of for scrap , and the sleepers loaded, some for disposal and some for use at Grosmont MPD.
Tues 4th Jan 2017 –
Rod, John & Nick in attendance today.
The final bolting up work and welding work, that was necessary to complete the north end hopper, was done today.
As the wagon is now over the pit, this also enabled the central gusset plate in the south end hopper to be patch welded. Thus the centre door can now be replaced to enable easier access to the interior.(Job for Thursday?) There is far less work to do at this end.
The pit access also allowed outstanding chassis and new metal paintwork to be progressed.
John gave this machine some TLC today, cleaning out the compacted swarf, then oiling & adjusting. There is a bit more to do next week, but John reckons it is now in good condition and not fit for condemning as has previously been voiced.
Thurs 5th Jan –
Bryan, Richard K and Nick in attendance today.
Bryan burning off failed paint on plough 18 prow.
Richard has sorted out the Webasto heater on the regulator and found a snazzy operating switch for it to replace the temporary one. The starter switch has had a lose terminal tightened and it works fine now. The stuck engine air filter warning light has been traced to a faulty connection.
Unfortunately, the wiring disturbance doing all this work, has caused the CCTV screen to sulk. It flashes the on blue light occasionally Steve- Richard has now handed this back to you for your expert attention).
We decided to wait until the Seacow is back over a solid floor before attempting to lift and fit the south end centre doors. Work now planned for a fortnight today. (19th).
I cleaned up all the recent welding work, ground off surface corrosion in the south end door rotation critical areas and applied primer, undercoat and gloss to various parts.
Peter Robinson of the Grosmont station group visited us today. He measured up Plough 18 as he si to make some new skylights and went away with surplus paint stocks – all the household type colours and stuff we don’t use on wagons and plant. A fair swop!
Sun 8th Jan – Gerry and Duncan
After a brief talk this morning over tea we decided to recover blocks off 214 so Duncan ground off the welds and we knocked them off easy. After cleaning up we tried new bush’s in and they fitted good.
Skates and short feeler were then taken off 250 and worn blocks cut off both sides. We then cleaned up the frames and positioned replacement blocks on one side and lined up ready for welding next week
Tues 10th Jan –
Due to widespread post Christmas pestilence only Steve and Nick in today. Apologies received from John, Richard and Rod. Get well soon gents, we are missing you all.
Steve spent the day cutting to fit bits of new metal for the side gusset plates in the south end of the Seacow. Now ready for Rod to weld next week.
I completed all the outstanding paintwork on the north end of the wagon.
I also had a good look around for bushes, as promised. I found a bag of steel ones, which to my untrained eye, look the same length and diameter as the tamper specimens. I tried one over the feeler rods on the bench, and in the fitting alongside. Looks like a better fit than the original ones!
Anyway, probably not what you want but I’ve left one with the feeler rods on the bench, for interest.
Thurs 12th Jan –
Spent most of the day today facilitating the lighting electricians. The new west side lights ( all seven of them ) are now up and running. Bring your sun glasses, as it is very bright, especially at the back of the shed, now.
They also looked at the failed (numbers 2-4 ) lights on the east side. One burnt out and two others have more than one fault. All three tried with a known working bulb to no avail. The electrician recommends that the most cost effective solution is to replace the innards of them with an LED type light. This is being priced up for Park Street.
When I was not required to help dismantle and assemble scaffolding towers, or pinch bar wagons etc, outstanding chassis painting on the south end of the seacow was completed.
I have also today tackled the next section of the west side of tamper 73250 – the usual corrosion treatment, filling, priming and undercoating and once again singing the praises of the new staging.
Sun 15th Jan –
Just Gerry working on tamper today.
I have had a go at fixing lining trolley lock air valve. A bit of a pigs ear to do, it would be better if they were spaced out as there is very little space to work. It took all day trying to work with mirror but in the end had to take the panel off to get to it. I have put tape on to stop latch from moving with a tie rap on as well. I found two spare valves but one had a push button and the other switched both ways.
I did not have time to test so will do that next week,
Tues 17th –
Thurs 19th –
Sun 22nd – Duncan, Alistair and Gerry working on tamper today.
Duncan and Alistair spent some time first preparing the job and setting up lights.
After coffee the feeler rod fixing blocks were set up on the left side and welded on before dinner. After dinner work started on the on the right side while the skate was attached to the left side. At 4pm after welding was complete we started to fit to fit the skate on the right side but the bolts did not line up. It was found that the outside block was too far out so it was then removed to find two shims in which were removed and the block re-welded back on, the skate then fitted OK. A few minor jobs need doing next week to finish off but the machine an now be moved if needed.
Tues 24th –
Rear camera system power supply.
After reconnecting the 24v-12v power supply convertor the camera system failed to operate.
This was today diagnosed as a faulty power supply unit. The 12v output was unstable and fluctuated wildly.
Traced to a failed joint between east side flexible and the bend to the scilencer. Proposed action is to split the joint and pack with exhaust jointing paste.
Engine bay cover gas struts.
Broken top mounting stud drilled out, tapped and new bolt fitted.
Second damper fitted.
Cab end engine bay cover now opens and closes with ease. Recommend that the engine bay cover is still opened and closed by two people, as with one person the cover does distort putting a twist on the hinge.
Gas struts for the battery bay cover to be ordered and fitted.
Desk engine and system warning lights.
Intermittent Air filter warning light illumination has yet to be resolved.
In addition the reliability of the oil pressure and engine overheat warning lights should also be questioned as these are never tested or illuminated during normal operation of the vehicle.
John’s response to Richard –
Sun 29th Jan – GerryFeeler rod repair now complete and tested OK. Left side is a bit stiff but after greasing started to loosen up.
Lining trolley lock also tested OK.
Only jobs now outstanding is cross level re-calibration and heater exhaust pipe to repair.
Tues 31st Jan –
Richards report.- Engine.
Top cover removed.
Compacted dust debris removed from the base of cylinder cooling fins. Engine overheat sensor located on the west side center cylinder injector mounting. Over heat warning bulb tested ok. Circuit yet to test by earthing the sensor terminal.
Further investigation of the intermittent blocked air filter warning light revealed a disconnected wire hanging down the side of the engine. When earthed this wire illuminates the blocked air filter warning light. When connected back onto the intake vacuum sensor the light is permanently illuminated. Vacuum sensor removed for testing and the disconnected wire secured.
Rear brush drive.
Drive chain adjusted. Intermittent failure to engage drive to be investigated this Thursday during functional test in the yard.
Battery bay cover failed gas struts removed in anticipation of the new gas strut delivery.
Camera power supply.
Steve tried the unit again with confirmed PSU case earth. Output voltage remained unstable, confirming internal component failure.
Thurs 2nd Feb –
Removed, successfully tested and refitted both sensors. West side sensor operates at a lower vacuum. Cab to sensor wire secured and left disconnected to test for vibration earths faults during next Thursdays site work. If Thursdays test is ok sensors will be wired back into circuit.
Traction circuit electrical fault.
24v supply to the main traction switch missing. Yellow supply wire traced back via a relay to a white wire that had become detached from the main ignition switch. Reconnected and traction now available.
Brush motor chain drive tested under power ok.
Camera power supply.
Redundant radio supply unit removed and tested for suitability. Unfortunately it is not suitable.
Hydraulic fluid leak.
While greasing Bryan noticed a pool of fluid on the floor under the cab front. This was traced to one of the main traction system cab gauge connections. Gauge released from the desk and connection tightened.
Main air tank reservoirs.
As these are soon due inspection, Nick confirmed that there is enough space to extract these under the side ploughs.
Tues 7th Feb –